help!!
Experiment #7 Pre-Lab Assignment Hint: for full credit, show all calculations, use proper sig. figs. and units Name Stoichiometric Analysis of a Mixture Lab Instructor Lab time Thurs. Tues. Lab Day Mon. An unknown mixture containing KCIO) was added to an ignition tube weighing 41.860 g. The tube and mixture was then found to weigh 43.772 q. After the first heating, the tube and its contents weighed 43.621 g. After the second and third heatings, it weighed 43.332 g. Answer the following questions and show your calculations clearly. 1. a. What mass of oxygen was liberated from the unknown mixture? b. What mass of KCIO3 was contained in the original mixture? 2 KCIOs (s) 2 KCI (s) + 3 O2 (g) What was the percent KCIO) in the original mixture? c. One should always wear safety goggles when working in a chemistry laboratory. Why is it especially important during this experiment? Be specific. 2.
3. What is the purpose of reheating the mixture several times until a constant mass is obtained? 4. The preparation of oxygen in this experiment is "catalysed" by MnO2. What is a catalyst? that incronses e rale ePa chemical sbslana rencten wlout toela underjou ypeinennt Chemiel chen 5. Explain why the rubber tubing must be removed from the water trough before allowing the test tube to cool. +0 prevent woter Unton tune as coolS fram Deng Snohad
Answer #1

2KClO3 --->2KCl +3O2

Mass of ignition tube, M1 = 41.860 g

Mass of ignition tube+mixture, M2 = 43.772 g

Initial Mass of mixture containing KClO3 M = 43.772 g - 41.860 g = 1.912 g

Final mass of mixture after heating, M4 = 43.332 g - 41.860 g = 1.472 g

Mass of oxygen liberated from the unknown mixture = 1.912 g - 1.472 g = 0.44 g

Molar mass of oxygen = 32 g/mol

Moles of oxygen liberated = Mass/molar mass = 0.44 g/32 g/mol = 0.01375 moles

Molar ratio of oxygen and KClO3 = 3:2

3 moles of oxygen ---> 2 moles of KClO3

0.01375 moles ---> == 0.01375 x 2/3 = 0.0275/3 = 0.00917 moles

Molar mass of KClO3 = 122.51 g/mol

Mass of KClO3 = moles x molar mass = 0.00917 mol x 122.51 g/mol = 1.123 g

Percent of KClO3 in the mixture = (Practical mass/initial mass) x 100 = (1.123 g/1.912 g) x 100 = 58.73 %

2) Goggles has to be used in this experiment because the decomposition of potassium chlorate is an exothermic reaction and liberate large amount of heat, flame and smoke.

KClO3 is a strong oxidising agent and should be cautious. It should be away from combustible materials.

3) The purpose of reheating the mixture several times until the constant mass is obtained in order to confirm the decomposition of KClO3 present in the initial mixture.

Constant mass indicates that all the KClO3 present in the initial mixture is decomposed to the KCl and oxygen. If not reheated some amount of KClO3 may retain and the accurate value of percent of KClO3 in the mixture cannot be obtained.